# CLCE: An Approach to Refining Cross-Entropy and Contrastive Learning for Optimized Learning Fusion

### v1

@@ 4.2 Transfer Learning
We now assess the transfer learning performance of our proposed CLCE. Here, adhering to the widely accepted paradigm for achieving state-of-the-art results, models are initialized with publicly-available weights from pretraining on ImageNet-21k (Deng et al., 2009) since they are state-of-the-art, and are fine-tuned on smaller datasets using our new loss function. We leverage 8 datasets: CIFAR-100 (Krizhevsky et al., 2009), CUB-200-2011 (Wah et al., 2011), Caltech-256 (Griffin et al., 2007), Oxford 102 Flowers (Nilsback and Zisserman, 2008), Oxford-IIIT Pets (Parkhi et al., 2012), iNaturalist 2017 (Horn et al., 2018), Places365 (Zhou et al., 2014), and ImageNet-1k (Deng et al., 2009). We adhere to official train/test splits and report mean Top-1 test accuracy over three different initializations.
@@
Tab. 2 presents the results of transfer learning, which offers further evidence of the effectiveness of our proposed CLCE approach beyond few-shot scenarios. When applied to four state-of-the-art image models, including BEiT-3, ResNet-101, ViT-B and MAE, our proposed CLCE approach consistently surpasses other methods, including the standard CE, H-SCL (Jiang et al., 2022) and the same weighted combination of CE and SupCon loss as CLCE. A paired t-test confirms these improvements as statistically significant ($p<0.05$). While the increase in performance with BEiT-3-CLCE over the BEiT-3-CE baseline is modest in some cases, such as the rise from $98.00\%$ (BEiT-3-CE) to $98.93\%$ (BEiT-3-CLCE) on CUB-200, it shows significant enhancements in challenging datasets with a higher level of class diversity. A notable example is iNaturalist2017, which has 5089 different classes, where CLCE leads to a marked improvement in accuracy from 72.31% to 75.72%. This substantial increase suggests that CLCE’s benefits are more pronounced in more varied datasets. In the case of ImageNet-1k, accuracy increased from $85.40\%$ (BEiT-3-CE) to $86.14\%$ (BEiT-3-CLCE), setting a new state-of-the-art for base models (88 million parameters)  [^1] https://paperswithcode.com/sota/image-classification-on-imagenet. We observe similar improvements in other transformer-based models, such as ViT and MAE. The use of CLCE in fine-tuning ResNet-101 also resulted in significant performance gains, particularly in the Caltech-256 dataset. Here, the model’s accuracy increases from $81.38\%$ (ResNet-101-CE) to $85.05\%$ (ResNet-101-CLCE). Compared to ResNet-101-CE, there has been an average increase in accuracy of $1.83\%$ for ResNet-101-CLCE. Furthermore, H-SCL (Jiang et al., 2022) yields inferior results compared to CE, mirroring the result observed in few-shot scenarios. Overall, the consistent achievement of high accuracies across diverse datasets using models fine-tuned with CLCE, especially ResNet-101 and BEiT-3, underscores the effectiveness of CLCE in improving model performance. Remarkably, this is achieved without resorting to specialized architectures, extra data, or heightened computational requirements, thereby establishing CLCE as a powerful alternative to traditional CE.
@@
Figure 2: Evaluation of the impact of the $\lambda$ hyperparameter. Results on eight tested datasets with $\lambda$ values ranging from $\{0,0.1,0.3,0.5,0.7,0.9,1.0\}$. The numerical details for these figures are provided in the supplementary material.
@@
SupCon (Khosla et al., 2020) 64 80.31 69.05
SupCon (Khosla et al., 2020) 128 82.17 69.93
@@
Table 3: Impact of different batch size. Performance of BEiT-3 base model when trained on CIFAR-FS and iNat2017 datasets. “CE” denotes cross-entropy loss. “SupCon” denotes supervised contrastive learning loss. “CLCE” denotes our proposed joint loss.
4.3 Reducing Batch Size Dependency
We evaluate the effect of batch size on the performance, specifically comparing our CLCE approach with CE and SupCon (Khosla et al., 2020). The results, as detailed in Tab. 3, indicate that SupCon’s performance is sensitive to batch size variations, a limitation not observed with CE. Particularly, SupCon shows inferior performance compared to CE with the commonly used batch size of 64 on both tested datasets. Even when the batch size is increased to 128, SupCon continues to underperform relative to CE. In our experiments, SupCon generally needs a batch size exceeding 512 to outperform CE, a requirement that is impractical for most single-GPU setups. This scenario mirrors the results of H-SCL (Jiang et al., 2022) in the context of few-shot and transfer learning. In contrast, CLCE not only surpasses CE performance on the iNat2017 dataset with a 1.41% accuracy improvement with batch size of 64 but also demonstrates an even more performance gain of 3.52% in accuracy with batch size of 128. Thus, our CLCE approach significantly mitigates the dependency on large batch sizes typically associated with contrastive learning approaches like SupCon and H-SCL. The reduction in dependency on large batch sizes greatly enhances the adaptability and effectiveness of CLCE in diverse computational settings, such as environments with budget GPUs equipped with 12 GB of memory.
@@ 4.4 Optimizing $\lambda$: Bridging CE and LACLN
Table 4: Results on CIFAR-FS and iNat2017 when training BEiT-3 base model using ablated versions of our CLCE. “CE” denotes cross-entropy loss. “CL” refers to our proposed label-aware contrastive learning, and “HNM” refers to hard negative mining.
@@ 4.6 Embedding Quality Analysis
To validate the enhancements brought by the proposed approach, CLCE, we perform a thorough evaluation focusing on the geometric characteristics of the generated representation spaces. We hypothesize that our CLCE enhances the quality of embeddings, thereby sharpening class distinction and improving performance. To elaborate, we examine the CE embeddings and CLCE embeddings produced by the BEiT-3 base model. Specifically, we evaluate two key aspects: (1)Distributions of cosine similarities between image pairs. This assessment provides insights into how well the model differentiates between classes in the embedding space. (2)Visualization of the embedding space using the t-SNE algorithm (Van der Maaten and Hinton, 2008). This visualization allows us to observe the separation or clustering of data points belonging to different classes.
We present the pairwise cosine similarity distributions of CE and CLCE embeddings in Fig. 3. Specifically, we randomly select the “tulips” class from CIFAR-100 to compute cosine similarities for positive (same class) and negative pairs (different classes). Observations from these plots reveal that the CLCE embeddings demonstrate superior separation between classes and less overlap between positive and negative samples compared to CE.

### v2

@@ 4.2 Transfer Learning
We now assess the transfer learning performance of our proposed CLCE. Here, adhering to the widely accepted paradigm for achieving state-of-the-art results, models are initialized with publicly-available weights from pretraining on ImageNet-21k [7] since they are state-of-the-art, and are fine-tuned on smaller datasets using our new loss function. We leverage 8 datasets: CIFAR-100 [24], CUB-200-2011 [55], Caltech-256 [14], Oxford 102 Flowers [40], Oxford-IIIT Pets [43], iNaturalist 2017 [20], Places365 [70], and ImageNet-1k [7]. We adhere to official train/test splits and report mean Top-1 test accuracy over three different initializations.
@@
Tab. 2 presents the results of transfer learning, which offers further evidence of the effectiveness of our proposed CLCE approach beyond few-shot scenarios. When applied to four state-of-the-art image models, including BEiT-3, ResNet-101, ViT-B and MAE, our proposed CLCE approach consistently surpasses other methods, including the standard CE, H-SCL [22] and the same weighted combination of CE and SupCon loss as CLCE. A paired t-test confirms these improvements as statistically significant ($p<0.05$). While the increase in performance with BEiT-3-CLCE over the BEiT-3-CE baseline is modest in some cases, such as the rise from $98.00\%$ (BEiT-3-CE) to $98.93\%$ (BEiT-3-CLCE) on CUB-200, it shows significant enhancements in challenging datasets with a higher level of class diversity. A notable example is iNaturalist2017, which has 5089 different classes, where CLCE leads to a marked improvement in accuracy from 72.31% to 75.72%. This substantial increase suggests that CLCE’s benefits are more pronounced in more varied datasets. In the case of ImageNet-1k, accuracy increased from $85.40\%$ (BEiT-3-CE) to $86.14\%$ (BEiT-3-CLCE), setting a new state-of-the-art for base models (88 million parameters)  [^2] https://paperswithcode.com/sota/image-classification-on-imagenet. We observe similar improvements in other transformer-based models, such as ViT and MAE. The use of CLCE in fine-tuning ResNet-101 also resulted in significant performance gains, particularly in the Caltech-256 dataset. Here, the model’s accuracy increases from $81.38\%$ (ResNet-101-CE) to $85.05\%$ (ResNet-101-CLCE). Compared to ResNet-101-CE, there has been an average increase in accuracy of $1.83\%$ for ResNet-101-CLCE. Furthermore, H-SCL [22] yields inferior results compared to CE, mirroring the result observed in few-shot scenarios. Overall, the consistent achievement of high accuracies across diverse datasets using models fine-tuned with CLCE, especially ResNet-101 and BEiT-3, underscores the effectiveness of CLCE in improving model performance. Remarkably, this is achieved without resorting to specialized architectures, extra data, or heightened computational requirements, thereby establishing CLCE as a powerful alternative to traditional CE.
4.3 Reducing Batch Size Dependency
We evaluate the effect of batch size on the performance, specifically comparing our CLCE approach with CE and SupCon [23]. The results, as detailed in Tab. 3, indicate that SupCon’s performance is sensitive to batch size variations, a limitation not observed with CE. Particularly, SupCon shows inferior performance compared to CE with the commonly used batch size of 64 on both tested datasets. Even when the batch size is increased to 128, SupCon continues to underperform relative to CE. In our experiments, SupCon generally needs a batch size exceeding 512 to outperform CE, a requirement that is impractical for most single-GPU setups. This scenario mirrors the results of H-SCL [22] in the context of few-shot and transfer learning. In contrast, CLCE not only surpasses CE performance on the iNat2017 dataset with a 1.41% accuracy improvement with batch size of 64 but also demonstrates an even more performance gain of 3.52% in accuracy with batch size of 128. Thus, our CLCE approach significantly mitigates the dependency on large batch sizes typically associated with contrastive learning approaches like SupCon and H-SCL. The reduction in dependency on large batch sizes greatly enhances the adaptability and effectiveness of CLCE in diverse computational settings, such as environments with budget GPUs equipped with 12 GB of memory.
@@ 4.3 Reducing Batch Size Dependency
Figure 2: Evaluation of the impact of the $\lambda$ hyperparameter. Results on eight tested datasets with $\lambda$ values ranging from $\{0,0.1,0.3,0.5,0.7,0.9,1.0\}$. The numerical details for these figures are provided in the supplementary material [35].
@@
SupCon [23] 64 80.31 69.05
SupCon [23] 128 82.17 69.93
@@
Table 3: Impact of different batch size. Performance of BEiT-3 base model when trained on CIFAR-FS and iNat2017 datasets. “CE" denotes cross-entropy loss. “SupCon" denotes supervised contrastive learning loss. “CLCE" denotes our proposed joint loss.
@@ 4.4 Optimizing $\lambda$: Bridging CE and LACLN
Table 4: Results on CIFAR-FS and iNat2017 when training BEiT-3 base model using ablated versions of our CLCE. “CE" denotes cross-entropy loss. “CL" refers to our proposed label-aware contrastive learning, and “HNM" refers to hard negative mining.
@@ 4.5 Ablation Study
Figure 4: Plot of cosine similarity distribution across the “cloud” class from CIFAR-100. Blue represents similarities of positive samples, while orange represents similarities of negative samples.
@@ 4.6 Embedding Quality Analysis
To validate the enhancements brought by the proposed approach, CLCE, we perform a thorough evaluation focusing on the geometric characteristics of the generated representation spaces. We hypothesize that our CLCE enhances the quality of embeddings, thereby sharpening class distinction and improving performance. To elaborate, we examine the CE embeddings and CLCE embeddings produced by the BEiT-3 base model. Specifically, we evaluate two key aspects: (1) Distributions of cosine similarities between image pairs. This assessment provides insights into how well the model differentiates between classes in the embedding space. (2) Visualization of the embedding space using the t-SNE algorithm [53]. This visualization allows us to observe the separation or clustering of data points belonging to different classes. (3) We employ the Isotropy Score as defined by [38] to evaluate the quality of produced embeddings. The Isotropy Score measures the distribution of data in the embedding space and serves as a metric for the quality of the produced embeddings. Historically, isotropy has served as an evaluation metric for representation quality [1]. This is based on the premise that widely distributed representations across different classes in the embedding space facilitate better distinction between them.
We present the pairwise cosine similarity distributions of CE and CLCE embeddings in Figs. 3 and 4. Specifically, we randomly select the “tulips” and “cloud” classes from CIFAR-100 to compute cosine similarities for positive (same class) and negative pairs (different classes). Observations from these plots reveal that the CLCE embeddings demonstrate superior separation between classes and less overlap between positive and negative samples compared to CE.
@@
Formally, we calculate the quantitative Isotropy Score (IS) [38], which is defined as follows:
$IS(\mathcal{V})=\frac{max_{c\subset C}\sum_{v\subset V}\exp{(C^{T}V)}}{min_{c\subset C}\sum_{v\subset V}\exp{(C^{T}V)}}$ (4)
where $V$ is a set of vectors, $C$ is the set of all possible unit vectors (i.e., any $c$ so that $||c||=1$) in the embedding space. In practice, $C$ is approximated by the eigenvector set of $V^{T}V$ ($V$ are the stacked embeddings of $v$). The larger the IS value, the more isotropic an embedding space is (i.e., a perfectly isotropic space obtains an IS score of 1).
Model iNaturalist2017 Imagenet-1k Places365
BEiT3-CE 0.32 0.27 0.34
BEiT3-CLCE 0.98 0.92 0.93
Table 5: Comparison of Isotropy Score across three datasets for BEiT-3-CE and BEiT-3-CLCE. A higher value is better. A higher Isotropy Score indicates better isotropy and generalizability.
Tab. 5 demonstrates that the IS score for BEiT-3-CLCE is superior to that of BEiT-3-CE, confirming that CLCE produces a more isotropic semantic space. The BEiT-3-CE embeddings are more anisotropic, implying that BEiT-3-CLCE embeddings more distinctly separate the different classes.